The solution to this equation (3a-3)-3(4a+5)+6=0
Answer by farohw(175)


(3a-3)-3(4a+5)+6=0
3a-3-12a-15 + 6 = 0
-9a - 18 +6 = 0
-9a - 12 = 0
-9a = 12
Divide both sides by -9 to get a = -12/9
Reduce this to lowest terms:
a = -4/3 is your answer.